OpenStudy (anonymous):

Q6. Suppose that the New England Colonials baseball team is equally likely to win a game as not to win it. If 5 Colonials games are chosen at random, what is the probability that exactly 3 of those games are won by the Colonials?

OpenStudy (sohailiftikhar):

it means u have to find that probability in which they only lose 2 games out of 5

OpenStudy (phi):

Use the binomial distribution \[ \Pr(\text{ k wins out of n tries})= \left(\begin{matrix}n \\ k\end{matrix}\right) p^k (1-p)^{n-k}\] in your problem n=5 , k=3 p= 0.5 ( 50% chance to win) and 1-p is also 0.5